I have homework for Algebra. I tried everything I know how to do. Can you help by showing me how to do it? The problem is -.3x+y=.5 , .5x-.3y=1.9 I know the answer because it was given to me by the yeacher. I just have to show how to do it. Please help!!!!!
Answer by stanbon(75887)

-.3x+y=.5
.5x-.3y=1.9
---------------
Multiply thru each equation by 10 to get:
-3x + 10y = 5
5x - 3y = 19
-----
Multiply thru the top equation by 5
Multiply thru the bottom equation by 3
-----
-15x + 50y = 25
15x - 9y = 57
-------------------
Add and solve for "y":
41y = 82
y = 2
----
Solve for "x":
15x - 9y = 57
15x - 18 = 57
15x = 75
x = 5
==================
Cheers,
Stan H.
====================